### v1

Abstract
The current evaluation of Large Language Models (LLMs) predominantly relies on benchmarks focusing on their embedded knowledge by testing through multiple-choice questions (MCQs), a format inherently suited for automated evaluation. Our study extends this evaluation to explore LLMs’ pragmatic competence—a facet previously underexamined before the advent of sophisticated LLMs, specifically in the context of Korean. We employ two distinct evaluation setups: the conventional MCQ format, adapted for automatic evaluation, and Open-Ended Questions (OEQs), assessed by human experts, to examine LLMs’ narrative response capabilities without predefined options. Our findings reveal that GPT-4 excels, scoring 81.11 and 85.69 in the MCQ and OEQ setups, respectively, with HyperCLOVA X, an LLM optimized for Korean, closely following, especially in the OEQ setup, demonstrating a score of 81.56 with a marginal difference of 4.13 points compared to GPT-4. Furthermore, while few-shot learning strategies generally enhance LLM performance, Chain-of-Thought (CoT) prompting introduces a bias toward literal interpretations, hindering accurate pragmatic inference. Considering the growing expectation for LLMs to understand and produce language that aligns with human communicative norms, our findings emphasize the importance for advancing LLMs’ abilities to grasp and convey sophisticated meanings beyond mere literal interpretations.
Index Terms:
Large Language Model (LLM), LLM Evaluation, Pragmatics, Korean.
[…]
=-21pt
I Introduction
Research on LLMs has seen rapid advancements in recent years [1], [2], [3]. Notably, ChatGPT[4], a web-based service built on GPT-3.5[5], was released in November 2022 and exemplifies this remarkable technological advancement. It has demonstrated impressive capabilities in a broad spectrum of Natural Language Processing (NLP) tasks, ranging from traditional ones like sentiment analysis and translation [6], [7] to more demanding areas such as complex problem-solving and creative writing [8], [9]. This versatility of ChatGPT has not only drawn significant attention from NLP researchers but has also captivated the interest of the general public worldwide and numerous tech companies are rigorously entering into the development of their own LLMs [10], [11], [12].
The evaluation of LLMs is as crucial as their development, as it enables the measurement of their performance for the targeted tasks and ensures that these models align with the anticipated standards of capability [13]. Systematic evaluation allows uncovering both the strengths and weaknesses of LLMs, which makes further fine-tuning of the models possible. This cycle of development and evaluation is essential in advancing these models, contributing to the creation of more reliable and suitable LLMs across diverse domains of real-world applications [14], [15].
Benchmarks are serving a significant role in the current evaluation of LLMs [16], [17]. By providing task-specific datasets aligned with corresponding metrics, they establish standardized scenarios in which fine-tuned LLMs can be consistently evaluated. The primary advantage of these benchmarks lies in their ability to automate the evaluation process, allowing a fair comparison of LLMs trained by different developers using varied strategies. However, there are notable limitations in the current benchmarking approach: a predominant focus on aspects like reasoning, computation, and knowledge [18], [19], [20] with an emphasis on literal meaning, rather than implied meanings varying with context, which may not fully capture human-like language understanding. Additionally, many benchmarks rely on MCQs, a format that, while convenient for automated evaluation, does not fully evaluate the generative capacities of LLMs [21]. Furthermore, current benchmarks are showing clear tendency to English-centric evaluation, which results in the under-exploration of LLMs’ multilingual capacities [22], [23].
TABLE I: Variations in Pragmatic Interpretation Based on Context
Statement ”There’s pizza in the fridge.
Literal meaning Pizza is present inside the fridge.
Implicated meaning 1 You are allowed to eat the pizza.
Implicated meaning 2 I won’t cook dinner for you.
Pragmatics is a linguistic study dealing with understanding language beyond just the literal meanings of words. It involves interpreting both the explicit (literal) and implicit (nonliteral) aspects of language, heavily influenced by context. Grice, H. P. illuminates this by distinguishing between what is said (literal meaning) and what is implicated (nonliteral meaning), emphasizing the critical role of context in language comprehension [24]. As an illustration, consider Table I, which presents a scenario involving the statement ”There’s pizza in the fridge.” The literal interpretation is straightforward: pizza is inside the fridge. However, the implicated meanings vary with context. In the first scenario, imagine a friend visiting and expressing hunger; the statement might imply permission to eat the pizza. In the second scenario, consider a couple recovering from an argument. Here, the same statement might carry an undertone of reluctance to cook, reflecting the strained atmosphere. These examples illustrate that the pragmatic interpretation of a simple statement can vary significantly, transforming it into complex, context-dependent communication.
While earlier NLP models primarily focused on syntactic and semantic aspects of human language, with little emphasis on pragmatics, current LLMs necessitate a more comprehensive evaluation that extends beyond these traditional aspects [25], [26]. The enhanced performance of these models across diverse NLP tasks marks the significant need for evaluating their contextual language comprehension. Pragmatic understanding is especially crucial for conversational setups where AI assistants are expected to understand and respond in ways that meet human communicative needs [27], [28].
In this paper, we demonstrate a systematic evaluation of LLMs’ pragmatic competence for Korean by analyzing it through four Gricean maxims: quantity, quality, relation, and manner, which are essential for understanding conversational implicature. We parallelize our experiments in two different setups: MCQs to explore LLMs’ pragmatic abilities through automated evaluation and OEQs to assess their narrative generation capacities, thereby revealing biases depending on question types. Furthermore, we investigate the impact of few-shot learning [29] and CoT reasoning [30], in-context learning strategies known to enhance model performance without the need for extensive retraining. By focusing on Korean, we demonstrate LLMs’ capabilities in handling Korean and provide deeper insights into their pragmatic capabilities, including those specifically optimized for Korean. Our work aims to narrow the existing gap between the rapidly evolving capabilities of LLMs and the nuanced, human-level evaluation of language, ultimately contributing to the development of AI systems with a more sophisticated awareness of contextual nuances.
II Related Work
II-A Evaluation Landscape: Methods for Assessing LLMs
The demand for an automated evaluation of language models that combines speed and reliability has become increasingly apparent, aiming to replace the costly and time-consuming process of human assessments. To meet this need, a range of benchmarks have been developed, each designed to evaluate the varied capabilities of different language models with specific focuses [31], [32], [33]. However, as LLMs emerge and evolve, these existing benchmarks have gradually begun to lose their discriminative power, leading to the creation of new benchmarks equipped with more challenging datasets—a task that remains an ongoing process [34], [35], [36]. Today, those benchmarks stand as the primary indicators for claiming the superiority of newly released LLMs.
The qualitative evaluation of generated text is a crucial part of assessing LLMs, as they have evolved as generative language models [37]. Predominantly, current benchmarks are designed to assess the LLMs’ capabilities in language understanding, thus inheriting a notable limitation in a thorough measurement of their generative abilities [18], [19], [20]. In this context, the role of human evaluations becomes indispensable for bridging these existing gaps. By providing a detailed assessment, human evaluation is adept at revealing both the strengths and weaknesses of LLMs at a finer-grained level. Consequently, human evaluation serves to complement aspects that automated evaluations may overlook, and continues to retain a critical importance in the ongoing research of LLMs [38], [39].
Recently, there has been active research into using LLMs as tools for evaluation, a method known as LLM-based evaluation [40], [41], [42]. These evaluation frameworks have shown superior performance compared to other state-of-the-art automated evaluation models, but they still do not reach the level of expert groups and possess limitations such as bias towards long sentences and the text generated by LLMs [42], [43].
II-B Evaluating Pragmatic Competence of LLMs
There have been efforts to assess the pragmatic capabilities of LLMs across varied contexts. Pietro et al. [44] evaluated ChatGPT’s pragmatic ability in Italian using the Test for the Assessment of Pragmatic Abilities and Cognitive Substrates (APACS) [45] as a primary test set, which covers six categories—interview, description, narratives, figurative language 1, 2, and humor. Their findings indicate that, while ChatGPT broadly approximates the pragmatic comprehension capabilities of humans, it has shown a tendency towards over-informativeness. Further challenges are noted in computing text-based inferences, processing physical metaphors, and understanding jokes.
However, the methodology employed raises questions regarding the appropriateness of using human-targeted tests for AI systems. ChatGPT was observed to offer responses considered excessive when addressing questions with inherent human attributes, such as ”Do you live with your family?” or ”What is your house like?”, often by reasserting its AI identity. Viewing these responses as overly informative within this experimental setup overlooks the functional considerations of ChatGPT’s design as an AI assistant, which aims to adhere to AI ethics principles, including self-awareness [46]. Additionally, the lack of transparency concerning the test’s specifics, including the test details, the total number of questions, and the evaluation criteria, undermines the robustness of the experimental evidence.
Bojic et al. [47] evaluated the contextual understanding of implicature in LLMs according to Grice’s Cooperative Principle and its four maxims [24]. The study compared the performance of the selected LLMs—GPT-2, GPT-3, GPT-3.5, GPT-4 and Bard—with that of humans, which was tested in English. They reported that the performance of the LLMs surpassed the human average, with GPT4 showing the highest scores, outperforming all other models and human participants.
Yet, this study is limited in that the human participants were recruited from the engineering faculty of the University of Belgrade in Serbia, specifically those who passed an internally developed English test. This suggests that the evaluation may not reflect the pragmatic competence of native English speakers accurately. Furthermore, the limited number of test items—four per category, culminating in a total of twenty—is deemed insufficient for a representative evaluation of general pragmatic competence, both within each specific category and overall.
II-C Korean-Specific LLMs and its Evaluation
Recently, Korean IT companies have been at the forefront of developing Korean-specific LLMs. In August 2023, NAVER unveiled Clova X, an AI assistant service powered by HyperCLOVA X [48], an LLM specially optimized for Korean. This service is targeting to serve the unique demands of South Korea, aiming to fulfill local service needs. Furthermore, in January 2024, Kakao Brain released Honeybee [49], a Multimodal Large Language Model (MLLM), which integrates language and vision, as open source contributing to further LLM research.
Meanwhile, the development of Korean-specific open-source LLMs is being accelerated and centered around the Open Ko-LLM Leaderboard [50]. This leaderboard challenges LLMs with five benchmarks: four are Korean-adapted versions of existing English benchmarks (ARC [18], HellaSwag [51], MMLU [20], TruthfulQA [52]), and one is the independently developed Ko-CommonGen benchmark [53], which assesses generation capabilities in Korean. As of February 2024, the leaderboard has seen submissions of over thousand LLMs.
Separate from the leaderboard, KMMLU [54] emerges as another critical benchmark for assessing LLMs’ capabilities in Korean. Specifically designed to test performance across 45 diverse categories, it focuses on a broad spectrum of Korean-specific knowledge. This benchmark is anticipated to further facilitate the development of LLMs that accurately embody the region-specific knowledge of Korea.
While not targeting an LLM, Nam et al. [55] evaluated the pragmatic competence of an AI speaker in Korean, using Gricean maxims in a multi-turn dialogue setup. They found that the maxim of relation was the most frequently violated by the AI speaker. Yet, it is evident that research into the pragmatic abilities of Korean LLMs remains in its initial phases, highlighting both a lack of studies specialized for Korean and the need for more in-depth evaluations of LLMs’ pragmatic understanding.
nanummj
TABLE II: Example of Test Unit on the Maxim of Quality
Korean English Translation
Context &
Statement
철수가 바이올린을 연습하자 옆에서 듣던 영희가 말했다.
”우리집 강아지가 더 잘 한다.”
As Cheol-su was practicing the violin, Yeong-hee, who was listening
next to him, said,
”My dog plays it better.”
MCQ
다음 보기에서 위 발화가 갖는 가장 적절한 의미를 고르세요.
(1) 철수의 바이올린 연주가 형편없다.
(2) 영희가 기르는 강아지는 철수보다 바이올린을 잘 켠다.
(3) 철수의 바이올린 연주는 강아지도 감동시킬 만큼 훌륭하다.
(4) 철수는 고양이를 키우고 있다.
Choose the most appropriate meaning of the statement above from the
options below.
(1) Cheol-su’s violin performance is terrible.
(2) The dog raised by Yeong-hee plays the violin better than Cheol-su.
(3) Cheol-su’s violin performance is so excellent that it can even move
    a dog.
(4) Cheol-su is raising a cat.
OEQ
위 발화가 갖는 가장 적절한 의미를 서술하세요.
Describe the most appropriate meaning of the statement above.
III Methodology
Our study is conducted across three distinct experiments designed to assess the pragmatic competence of LLMs. We begin by evaluating LLMs’ ability to make pragmatic inferences through MCQs, then move to analyze their narrative responses to OEQs to probe their generative capabilities. Lastly, we explore LLMs’ adaptability within in-context learning scenarios, incorporating few-shot learning and CoT prompting.
To further elaborate on our methodology, this section is structured as follows: Section 3.1 introduces Gricean conversational maxims, which form the theoretical foundation for our study. Section 3.2 details the development of a pragmatic test set, crafted specifically to assess the LLMs’ abilities in line with these maxims. Section 3.3 discusses two distinct test setups for MCQs and OEQs, focusing on assessing LLMs’ understanding and response generation. Section 3.4 outlines in-context learning scenarios, detailing our approach to evaluate the effects of few-shot learning and CoT techniques on LLMs’ pragmatic capabilities.
III-A Gricean Conversational Maxims
In evaluating the pragmatic competence of LLMs, we utilize Grice’s theory of conversational implicature as the foundational framework. Implicature refers to the meanings that speakers imply but do not explicitly state, which listeners must deduce from contextual cues. This aspect is essential for effective communication as humans often rely on implied meanings rather than explicit statements in real-world conversations.
Grice delineated the cooperative principle as the foundation for rational conversation. This principle states that participants should make contributions that are appropriate for the current stage of the conversation, guided by its accepted purpose or direction [24]. It is further categorized into four conversational maxims:
  • The Maxim of Quantity: Make your contribution as informative as is required.
  • The Maxim of Quality: Try to make your contribution one that is true.
  • The Maxim of Relation: Ensure that all the information you provide is relevant to the current conversation.
  • The Maxim of Manner: Be perspicuous; Be brief and orderly, and avoid obscurity and ambiguity.
Conversational implicatures are often expressed by intentionally flouting Grice’s maxims, which leads to implicated meanings beyond the literal. A common way to flout the maxim of quantity is through tautology—statements that are redundant or always true, offering no information, such as, “Either the news is true, or it is not true.” Flouting the maxim of quality can involve making obviously false or exaggerated statements, exemplified by, “Sally ran faster than light.” Similarly, flouting the maxim of relation occurs when statements are unrelated to the ongoing conversation, for instance, mentioning the weather during a conversation about exam results—“The weather today is very nice, isn’t it?”—demonstrates this principle. An example of flouting the maxim of manner, which requires clear and orderly expressions, might be a response to a political inquiry at a press conference with, “I would not try to alter your viewpoint about it,” intentionally avoiding a direct answer to maintain ambiguity.
III-B Constructing Pragmatic Test Set
Guided by Grice’s maxims, we have constructed a test set consisting of 120 test units, with 30 units allocated to each maxim. Each test unit is designed to elicit a specific implicated meaning, facilitating an evaluation of LLMs’ capabilities in pragmatic inference. This set was developed collaboratively by four experts holding master’s degrees in linguistics or related fields. Table II showcases an example from our test set, which was originally crafted in Korean and is provided with its English translation. Each test unit consists of a context that sets the scene for the dialogue, a statement made by one of the participants, and a follow-up question that asks the expressed meaning of the statement.
This test set is designed to challenge LLMs across various conversational scenarios, assessing their ability to interpret implicature in a manner akin to human understanding. By analyzing LLMs’ responses to the test set, we aim to quantify the level of their pragmatic competence, offering insights into both their capabilities and limitations in processing human language. The complete test set is available at https://anonymous.4open.science/r/pragmatic_eval_korean-43D0/.
TABLE III: Experimental Setups for Assessing LLMs’ In-Context Learning Capabilities in the MCQ Test
Setup Few-Shot Examples CoT Prompting
0-shot (Base) 0 No
1-shot (Base) 1 No
4-shots (Base) 4 No
0-shot (CoT) 0 Yes
1-shot (CoT) 1 Yes
4-shots (CoT) 4 Yes
III-C Multiple-Choice Questions vs. Open-ended Questions
Table II illustrates the variety of question types accompanying each given context and statement within a test unit, categorized by their format into either MCQs or OEQs. MCQs require selecting the most appropriate meaning of the statement among the provided options. Conversely, OEQs involve narratively describing the most appropriate meaning of the statement, without preset options. LLMs are tested in both setups for analyzing differences in performance and behavioral patterns.
For MCQs, each question is accompanied by four options. We categorized these options to represent distinct types of interpretation: the correct answer that accurately reflects the implicated meaning within the context, a naive literal interpretation, an incorrect interpretation within context, and an incorrect interpretation out of context. A response is considered correct if the LLMs’ generated answer explicitly selects the option number corresponding to the correct interpretation.
For OEQs, we engage three independent assessors who meet the same qualifications as those involved in constructing the test set, to qualitatively assess the narrative responses from LLMs. The assessors rate the LLMs’ responses on a 1 to 5 scale as follows: A score of 5 signifies perfect contextual understanding and accurate interpretation of implicature. A score of 4 reflects an understanding of context with incomplete implicature interpretation. A score of 3 indicates a tendency to focus on the literal meaning, while a score of 2 suggests some misinterpretation of the literal meaning. Lastly, a score of 1 represents a comprehensive misunderstanding of both the context and the literal meaning. For ease of comparison with MCQ outcomes, we present these results as re-scaled scores ranging from 0 to 100. We ensured that informed consent was obtained from all human subjects involved in our study.
III-D In-Context Learning
Recent research highlights the effectiveness of in-context learning, enabling LLMs to rapidly adapt to unfamiliar tasks by leveraging their extensive pre-trained knowledge, without the need for prior training on those tasks [56], [57]. This study focuses on two specific in-context learning strategies: few-shot learning [29] and CoT prompting [30]. Detailed in Table III, our experiment utilizes the MCQ test format to investigate the impact of these strategies on the LLMs’ pragmatic competence, comparing six distinct experimental setups.
We delineate three few-shot learning scenarios: zero-shot, one-shot, and four-shots, each defined by the number of examples provided. The one-shot scenario includes a single test example unit representing the maxim of quality. In the four-shot scenario, four distinct example units are provided, each exemplifying a different maxim, thereby enriching the context provided to LLMs.
Our experiment further distinguishes between two setups based on the implementation of CoT prompting. The base setup omits the CoT approach, presenting solely the test examples and their respective answers. Conversely, the CoT setup includes detailed, step-by-step reasoning processes for the provided examples, encouraging LLMs to explicitly articulate its inferential process as shown in its few-shot examples.
nanummj Specifically, for the zero-shot scenario with CoT, we adopt the methodology of Kojima et al. [58], by appending the phrase “답: 순차적으로 생각해봅시다.” (translated as “Ans- wer: Let’s think step by step.”) at the end of the question, guiding the model towards a structured inferential approach.
IV Experiment
IV-A Experimental Setup
Model Selection
Our study considers five distinct LLMs to compare: GPT-3.5-turbo [59] (henceforth GPT-3.5), GPT-4 [60], Gemini-Pro [61], HyperCLOVA X [48], and LDCC-Solar-10.7B [62] (hereinafter LDCC-Solar). GPT-3.5 and GPT-4 are developed by OpenAI, and Gemini-Pro by Google DeepMind. These models are multilingual LLMs, capable of processing Korean among other languages. In contrast, HyperCLOVA X is developed by NAVER, a leading Korean IT company, and LDCC-Solar is the top-performing open-source model on the Open Ko-LLM Leaderboard as of February 2024, instruction-tuned by Lotte Data Communication using Solar [63] as the foundation model. Both HyperCLOVA X and LDCC-Solar are LLMs specifically optimized for processing Korean.
Hyperparameter Setting
For each LLM, we uniformly configured the hyperparameter settings to ensure a fair comparison across the board. Specifically, the maximum output length was set to 512 tokens, and the temperature parameter was adjusted to 0.7. For HyperCLOVA X, which was not accessible via API, we manually retrieved each response from its website, initiating a new session for each interaction to maintain consistency with the other LLMs.
Performance Report
To report the performance of each LLM, we generated three responses from each model for each test unit. In the case of the MCQ tests, we varied the order of the provided options to mitigate potential bias towards any specific answer position. We then report the averaged scores from three trials to ensure the reliability of our results, taking into account the variability in the outputs of the LLMs.
TABLE IV: Scores on MCQ Test Across four Gricean Maxims
Quantity Quality Relation Manner Avg.
GPT-3.5 36.67 50.00 28.89 44.44 40.00
[…]
Gemini-Pro 62.22 75.56 53.33 47.78 59.72
[…]
LDCC-Solar 57.78 57.78 36.67 45.56 49.44
IV-B Result of the MCQ Test
Analysis of LLM Performance
Table IV illustrates the performance of each LLM on the MCQ test across the four Gricean maxims, along with the overall average score. GPT-4 leads with an impressive average score of 81.11, significantly outperforming all compared LLMs. Following closely are HyperCLOVA X and Gemini-Pro, with scores of 67.5 and 59.72, respectively, showcasing their proficiency in understanding conversational implicature.
Notably, LDCC-Solar, despite possessing nearly half the parameter size of GPT-3.5 (10.7 billion compared to GPT-3.5-turbo’s 20 billion) [64], manages to exceed GPT-3.5’s performance by 9.44 points. This highlights the effectiveness of LLMs specialized for Korean, suggesting that a larger parameter size does not necessarily guarantee enhanced performance.
Conversely, the significant performance gap between GPT-3.5 and GPT-4, which boasts 1.7 trillion parameters, underscores the continued importance of parameter scale. This difference serves to emphasize that, while specialized training is crucial, the scale of parameters remains a critical factor in facilitating a model’s capabilities, particularly in tasks requiring nuanced pragmatic inference.
In the evaluation of individual maxims, LLMs consistently score higher on the maxim of quality, while they tend to receive lower scores for the maxims of relation and manner. This pattern suggests that, within the MCQ setup, the presence of options based on literal interpretations may inadvertently make it more challenging to choose answers that align with the appropriate implicature, particularly for the maxims of relation and manner. The reason behind this is twofold: For the maxim of quality, literal interpretations often lack semantic plausibility, making them easier for LLMs to rule out. In contrast, for the contexts governed by relation and manner, options with literal meanings can be often semantically valid, complicating the selection process.
Notably, HyperCLOVA X’s superior performance in the maxim of quality, surpassing even GPT-4, can likely be attributed to its language-specific training. This implies the advantage of developing a model with a comprehensive Korean dataset, which enables it to achieve a deeper understanding of intricate linguistic nuances.
Figure 1: MCQ Option Choice Distribution by each LLM
LLMs’ Answer Selection Examined
Figure 1 demonstrates the distribution of option types selected by the LLMs during the MCQ test, revealing a noticeable preference across all models for literal interpretations among wrong answer options. This tendency underscores the challenge posed by a bias towards literal meaning, which obstructs effective pragmatic inference. Specifically, GPT-3.5 shows a pronounced preference for literal interpretations, selecting them 53.6% of the time, compared to only 40% for correct pragmatic interpretations.
The analysis further reveals that, even when selecting incorrect answers, all LLMs are more inclined to choose options that are incorrect within the given context rather than options that are out of context. This behavior suggests that the models do engage with the context in their decision-making, albeit not always successfully leading to the correct implicature.
LDCC-Solar, in particular, shows a tendency to select options not provided in the prompt—such as ‘5’ or ‘d’, which it generates on its own, at a rate of 3.9%—underscoring a significant challenge in adhering to the given choices and further deviating from accurate pragmatic inference. Additionally, LDCC-Solar’s responses often include ‘garbage text’, irrelevant material that detracts from the response quality, pointing to challenges that go beyond understanding pragmatic cues.
IV-C Result of the OEQ Test
TABLE V: Scores on OEQ Test Across four Gricean Maxims
Quantity Quality Relation Manner Avg.
GPT-3.5 61.25 49.75 67.50 64.75 60.81
[…]
Gemini-Pro 75.00 64.75 77.00 68.50 71.31
HyperClova X 81.50 83.50 88.00 73.25 81.56
LDCC-Solar 62.25 54.00 62.50 49.25 57.00
Performance Evaluation from MCQs to OEQs
In Table V showcasing the results of the OEQ test, GPT-4 maintains its leading performance, with HyperCLOVA X and Gemini-Pro closely following, consistent with the MCQ test findings. However, while the performance gap between GPT-4 and HyperCLOVA X was 13.61 in the MCQ test, it has significantly narrowed to 4.13 in the OEQ test, indicating HyperCLOVA X’s enhanced capability in the narrative setup.
Conversely, LDCC-Solar, which outperformed GPT-3.5 in the MCQ test by 9.44, falls behind GPT-3.5 by 3.81 in the OEQ test, suggesting that while LDCC-Solar excels in option selection, GPT-3.5 demonstrates superior generative abilities.
These discrepancies underscore a critical consideration for LLM development, highlighting that the MCQ framework, while popular for benchmarking, may not fully capture the essence of LLMs’ generative capabilities. Given that LLMs often operate in conversational setups without predefined options in real-world scenarios, the significance of qualitative evaluation in assessing LLM narrative generation capabilities becomes evident.
For all maxims, GPT-4 consistently achieves the highest scores, closely followed by HyperCLOVA X. Similar to the MCQ test, the maxim of manner yields lower scores overall, yet notably, the maxim of relation receives the highest scores in the OEQ test, with the maxim of quality scoring relatively lower, which are contrastive to the MCQ results. This shift likely suggests that the narrative format of OEQs, devoid of predefined options, simplifies the task of adhering to generating correct answers at the maxim of relation for LLMs, reducing its complexity. Conversely, this format appears to diminish the advantages previously observed for the maxim of quality in the MCQ setup.
Interestingly, LDCC-Solar exhibits a unique behavior by initially generating options akin to those in the MCQ format and then selecting one, even when asked to respond narratively. This behavior likely results from overfitting to the MCQ format, which predominates the benchmarks used in the Open Ko-LLM Leaderboard. Such a strategy, while innovative, may not always align with the expectations for narrative answers and could reflect a limitation in the model’s adaptability to varied question formats. Additionally, as observed in the MCQ test, LDCC-Solar’s responses often include irrelevant text that diminishes the overall quality of its responses.
Analyzing Score Distributions of LLMs
Figure 2 presents box plots that illustrate the score distributions of each LLM in the OEQ test. In these plots, GPT-4 and HyperCLOVA X exhibit the same median (Q2) and 75th percentile (Q3) values, indicating comparable performance at the median and upper quartiles. However, at the 25th percentile (Q1), representing the lowest 25% of scores, GPT-4 demonstrates superior performance with a score of 75 compared to HyperCLOVA X’s 66.75. This suggests that GPT-4 maintains a higher level of quality for the lower-performing test units, indicating greater overall stability in its responses.
Gemini-Pro displays a broader interquartile range between Q1 and Q3, indicating less consistency in its performance compared to GPT-4 and HyperCLOVA X. This wider range suggests variability in the quality of Gemini-Pro’s responses across different test units.
Similarly, GPT-3.5 and LDCC-Solar show no differences in their Q2 and Q3 values. However, the Q1 score for LDCC-Solar is markedly lower at 25 compared to 39.62 for GPT-3.5. This highlights that the lower-quality responses from LDCC-Solar contribute significantly to the reduced consistency in response quality.
IV-D Results of In-Context Learning
Figure 2: Box Plots of OEQ Score Distributions by each LLM
Table VI showcases the impact of employing two in-context learning techniques in the MCQ setup: few-shot learning and CoT prompting. Reiterating our findings from the MCQ test, GPT-4 continues to outperform all other LLMs across different setups, with HyperCLOVA X and Gemini-Pro following closely. The application of the few-shot learning strategy leads to incremental improvements in performance for most LLMs in both setups with and without CoT prompting.
Notably, HyperCLOVA X exhibits a remarkable improvement by few-shot learning in the base setup, achieving a higher increase compared to other LLMs; it gains 10.56 points moving from 0-shot to 1-shot learning, and an additional 6.08 points transitioning from 1-shot to 4-shots, resulting in a score of 84.14, which approaches GPT-4’s leading 89.17. Such impressive improvements indicate HyperCLOVA X’s underlying capacities in pragmatic competence and highlight its effectiveness as a Korean-specific LLM, which has evidently benefited from comprehensive pre-training on a diverse Korean text corpus.
In contrast to the advantages observed with the few-shot technique, applying CoT prompting appears to diminish the performance of LLMs compared to their base setups without CoT. Notably, the CoT approach tends to introduce a bias towards literal interpretations. This tendency is particularly evident in the 4-shot setup, where LLMs frequently select multiple options, often including those of literal interpretations.
Considering that the contexts provided in the evaluation were limited to 1-2 sentences, with information crucial for pragmatic inferences not explicitly delineated at the semantic level, the CoT method arguably obstructs the inherent capabilities of LLMs for pragmatic inference. This highlights a limitation of CoT, especially in scenarios where pragmatic cues are subtly embedded below the semantic level and require nuanced interpretation. While CoT has demonstrated potential in enhancing logical thinking and problem-solving in contexts with explicit statements, its effectiveness for pragmatic inference appears contingent on the depth and type of contextual information provided.
TABLE VI: Scores on MCQ Test Under Few-Shot Learning Conditions with and without Chain of Thought prompting
0-shot 1-shot 4-shots
base CoT base CoT base CoT
GPT-3.5 40.00 35.56 44.72 40.56 56.94 47.50
[…]
Gemini-Pro 59.72 59.17 63.06 61.94 66.39 62.78
HyperClova X 67.50 66.67 78.06 62.50 84.14 71.94
LDCC-Solar 49.44 52.50 59.44 0.00 0.00 0.00
nanummj Differently from other LLMs, LDCC-Solar exhibits a slight increase in score for the 0-shot CoT setup, indicating some initial advantages of this strategy. However, its performance drastically declines in the 1-shot with CoT and both 4-shot setups, with and without CoT, where it completely fails to generate the required answers, resulting in scores of 0. This failure appears to stem from LDCC-Solar’s tendency to reiterate the given prompt in its responses, which suggests a limitation in its processing capabilities when faced with increased complexity or specificity in the tasks. In most responses, it either repeats the prompts without adding any substantive content or produces irrelevant single words, such as ’철수’ (Cheol-su, a person’s name used in prompts), or meaningless fragments like ’제’ or ’먼’. These responses highlight the model’s difficulty in moving beyond a straightforward single-shot example, possibly reflecting the limits of its capacity to handle nuanced or layered instructions.
TABLE VII: Examples of Korean Culture-Specific Test Units
ID Type of Maxim
Context and Statement in Korean
English Translation
27
Quantity
붕어빵 가게 앞을 지나가던 철수가 영희에게
현금이 있는지 물었고 영희는 다음과 같이 말했다.
‘5만원짜리 밖에 없어’.
As Chul-su was passing by a 붕어빵(bung-eo-bbang)
street stall, he asked Yeong-hee if she had any cash.
Yeong-hee replied,
‘I only have a 50,000 won bill’.
83
Relation
철수 집에 놀러 간 영희는 주방에 많은 귤이 쌓여
있는 것을 보고 귤이 왜 이렇게 많은지 물었고
철수는 다음과 같이 말했다.
‘우리 작은 아버지께서 제주도에 사셔’.
When Yeong-hee visited Chul-su’s house, she saw
many tangerines piled up in the kitchen and asked
why there were so many. Chul-su replied,
‘My uncle lives on Jeju Island’.
V Case Study: In-Depth Analysis of LLM Responses to OEQs
V-A A Closer Look at Strengths and Weaknesses
In our analysis, we uncovered distinct characteristics within the responses of LLMs. Most notably, GPT-4, achieving the highest overall score in our evaluations, consistently delivers a singular, unambiguous answer rather than multiple potential interpretations. This decisiveness stands as a substantial strength, streamlining the decision-making process for end-users by eliminating the need to filter through various possible answers. Moreover, GPT-4’s strategy of providing precise responses, and avoiding the use of uncertain expressions like ‘-으로 예상할 수 있다’ (it can be speculated that) or ‘-일 수 있다’ (may be)—expressions that are frequently encountered in GPT-3.5 and occasionally in responses from Gemini and HyperCLOVA X—further contributes to the reliability of its answers.
HyperCLOVA X demonstrates its strengths by providing detailed explanations of its reasoning process, significantly facilitating a deeper comprehension for users. Conversely, Gemini-Pro reveals a critical limitation by occasionally offering terse responses without elucidation. An example of this is returning a single-word answer, ‘반어법’ (irony), without offering a supplementary interpretation.
V-B Analyzing Error Patterns in LLM Responses
Our examination of LLM responses has identified error patterns that highlight common interpretive challenges. The most prevalent error arises from the LLMs’ struggle to accurately interpret implicatures despite understanding the context, followed by errors due to purely literal interpretations, especially notable in scenarios related to the maxim of manner. This observation contrasts with our findings from the MCQ answer choice distribution analysis (cf. Section 4.2.2), where errors based on literal interpretations were more dominant. This discrepancy implies that the MCQ format, with its predefined options, may have encouraged LLMs to prefer literal meanings, complicating their ability to achieve accurate pragmatic inferences.
We also found that LLMs struggle with generating the correct interpretations when confronted with synonyms or phonetically similar words. A notable case in our test units involved the word ‘사기’ (pronounced as ‘sagi’), which ambiguously means both ‘to buy something’ and ‘a fraud’. Given that ‘사기’ is the nominalized form of ‘사다’ (‘sada’), meaning ‘to buy’, it was expected to be understood in the context of purchasing. Yet, among the LLMs evaluated, only HyperCLOVA X succeeded in accurately interpreting ‘사기’ as related to buying, whereas the others erroneously took it to mean fraud.
Furthermore, we observed errors associated with formatting in the responses. For example, Gemini-Pro occasionally included markdown formatting symbols into its output directly, like **표현** (translated as ‘expression’), with the intention of emphasizing the text. However, these markers appeared literally in the responses, resulting in inaccurately formatted and potentially confusing answers.
V-C Responses to Cultural-specific questions in Korean
Through our examination, we have observed that certain test units are deeply rooted in Korean culture and significantly influence the responses of LLMs. To exemplify this, we have chosen two prototypical instances, detailed in Table VII.
The first example features 붕어빵 (’bung-eo-bbang’), a beloved traditional Korean street food made by pouring batter into fish-shaped molds and filling them with sweet red bean paste, especially popular during the winter season. Typically sold at street stalls, bung-eo-bbang is expected to be purchased with cash, as these stalls often lack card reading facilities. The cost of bung-eo-bbang is approximately 1,000 won for a set of three. In the example provided, Yeong-hee mentions that she only has a 50,000 won bill. This statement can be interpreted as an expression of her reluctance to buy bung-eo-bbang, not just due to the inconvenience it would cause the stall owner in providing change, but also because of the impracticality for Yeong-hee to manage the received smaller bills. Thus, the statement ‘I only have a 50,000 won bill’ subtly conveys her unwillingness to make a purchase under these circumstances. While several LLMs, including GPT-4, struggled to accurately infer the implied meaning of Yeong-hee’s statement, HyperCLOVA X demonstrated a correct understanding, showcasing its ability to adjust responses based on the specific cultural context prevalent in Korea.
In the context of the second example, Jeju Island is renowned for its regional specialty, the Jeju-Tangerine, and it is a common practice among residents to send Jeju-Tangerines as gifts. Considering this context, the statement ’My uncle lives on Jeju Island’ can be implicitly understood to mean ’We have a lot of tangerines because my uncle, who lives on Jeju Island, sent them to us.’ Both GPT-4 and HyperCLOVA X excelled at figuring out the meaning behind the statement. Notably, even the LDCC-Solar, which generally scored lower on average across the board, achieved a high score in this specific case.
These findings align with the results of Son et al. [54], where HyperCLOVA X demonstrated its superior performance for Korea-specific knowledge compared to other LLMs, including GPT-4. This observation underscores the critical importance of developing LLMs that are capable of comprehending culture-specific contexts, which is essential not only for accurate general knowledge but also for higher-level processes, such as pragmatic inference.
VI Conclusion and Future Work
In this study, we address an under-explored aspect of LLM evaluation—the pragmatic evaluation of LLMs, with a specific focus on Korean. We developed a test set comprising 120 test units rooted in Gricean theory of conversational implicature, aiming to rigorously assess the pragmatic competence of LLMs, including those specialized for Korean.
Our findings reveal that GPT-4 surpasses all competing models in both MCQ and OEQ setups, with HyperCLOVA X and Gemini-Pro closely following. Notably, the comparative analysis between the MCQ and OEQ formats showcases HyperCLOVA X’s capability to significantly narrow the performance gap observed in the MCQ test. Similarly, LDCC-Solar, an open-source LLM, demonstrates superior performance to GPT-3.5 in the MCQ setup, yet falls short in the OEQ setup. These findings highlight the variability in LLM capabilities dependent on the question format, emphasizing the critical importance of incorporating an OEQ format for a holistic evaluation.
Furthermore, our exploration of in-context learning techniques yielded mixed results; few-shot learning generally enhanced LLM performance, whereas CoT prompting had a negative impact. We attribute this primarily to the nature of pragmatic expressions, which are often not explicitly stated at the semantic level, leading CoT to inadvertently emphasize literal interpretations.
While our study provides a comprehensive evaluation of LLMs’ pragmatic competence, we identify two primary areas for enhancement in our future work. Firstly, although the test set of 120 units—30 for each Gricean maxim—has yielded meaningful insights, this quantity remains modest in comparison to other benchmarks commonly utilized for LLM evaluation. Additionally, while focusing on Korean has unveiled significant findings, the multilingual capabilities of LLMs are yet to be fully explored. This highlights the necessity of extending our benchmark to other languages. However, such expansion requires more than mere translation; it demands careful consideration of cultural nuances associated with each language [54]. Viewing this study as a pilot, we aim to address these considerations in our future work, striving to develop a more robust and reliable framework for the multilingual evaluation of LLMs’ pragmatic competence.